### v1

# The Marcinkiewicz-Zygmund Property for Riemann Differences with Geometric Nodes

Abstract.
We study when a Riemann difference of order $n$ possesses the Marcinkiewicz-Zygmund (MZ) property: that is, whether the conditions $f(h)=o(h^{n-1})$ and $Df(h)=o(h^{n})$ imply $f(h)=o(h^{n})$. This implication is known to hold for some classical examples with geometric nodes, such as $\{0,1,q,\dots,q^{n-1}\}$ and $\{1,q,\dots,q^{n}\}$, leading to a conjecture that these are the only such Riemann differences with the MZ property (see [2]). However, this conjecture was disproved by the third-order example with nodes $\{-1,0,1,2\}$ ([5]), and we provide further counterexamples and a general classification here.
We establish a complete analytic criterion for the MZ property by developing a recurrence framework: we analyze when a function $R(h)$ satisfying $D(h)=R(qh)-AR(h)$, together with $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. We prove that this holds if and only if $A$ lies outside a critical modulus annulus determined by $q$ and $n$, covering both $|q|>1$ and $|q|<1$ cases. This leads to a complete characterization of all Riemann differences with geometric nodes that possess the MZ property, and provides a flexible analytic framework applicable to broader classes of generalized differences.
Key words and phrases:
Peano derivative, Riemann difference, MZ property, Taylor expansion
2020 Mathematics Subject Classification:
Primary 26A24; Secondary 26A27
1. Introduction
The concept of differentiation lies at the heart of analysis. To address limitations in smoothness and extend differentiation to broader function classes, various generalizations of the classical derivative have been developed. Among these, Riemann derivatives offer a discrete framework for computing higher-order derivatives directly from function values. These derivatives depend on a choice of interpolation nodes and associated coefficients, and their properties reveal subtle relationships between analytic and measure-theoretic structures.
Let $n\geq 1$ be fixed. For distinct real numbers $b_{1},b_{2},\dots,b_{n+1}$ (called nodes), the associated divided difference operator is the linear operator $\Delta_{n}$ defined on the space of real-valued functions $f$ by
$\Delta_{n}f=\sum_{k=1}^{n+1}\frac{f(b_{k})}{\prod_{j\neq k}(b_{k}-b_{j})}=\sum_{k=1}^{n+1}a_{k}f(b_{k}),$
where the weights $a_{k}=\frac{1}{\prod_{j\neq k}(b_{k}-b_{j})}$ depend only on the nodes. These weights coincide with the Lagrange interpolation coefficients corresponding to the nodes $b_{k}$.
The corresponding difference operator is defined by
$R_{h}f(c):=\sum_{k=1}^{n+1}a_{k}f(c+b_{k}h),$
so that the $n$-th Riemann derivative of $f$ at a point $c$ is given by
$Rf^{(n)}(c)=\lim_{h\to 0}\frac{n!R_{h}f(c)}{h^{n}},$
provided the limit exists. Notably, different choices of nodes give rise to different Riemann derivatives. For instance, the forward Riemann derivative corresponds to nodes $0,1,2,\dots,n$ and is given by
$Rf^{(n)}(c)=\lim_{h\to 0}\frac{\sum_{k=0}^{n}(-1)^{n-k}\binom{n}{k}f(c+kh)}{h^{n}},$
while a symmetric choice of nodes produces the symmetric Riemann derivative.
The weights $a_{k}$ associated to a Riemann derivative satisfy the system
(1.1) $\sum_{k=1}^{n+1}a_{k}b_{k}^{j}=\begin{cases}0&\text{for }j=0,1,\dots,n-1,\\ 1&\text{for }j=n.\end{cases}$
Thus, whenever the classical $n$-th derivative $f^{(n)}(c)$ exists, so does $Rf^{(n)}(c)$, and the two coincide. This positions Riemann derivatives as natural tools for discrete differentiation. However, the converse implication
$Rf^{(n)}(c)\Rightarrow f^{(n)}(c)$
does not hold in general. A simple counterexample is the Schwartz derivative:
$f_{s}(c)=\lim_{h\to 0}\frac{f(c+h)-f(c-h)}{2h},$
which exists at $c=0$ for $f(x)=|x|$ but $f$ fails to be differentiable at 00.
Peano observed that the existence of a polynomial $p(x)$ satisfying
$f(c+h)-p(c+h)=o(h^{n})$
does not imply the existence of $f^{(n)}(c)$. Functions satisfying this approximation property are said to be $n$-times Peano differentiable at $c$, and the coefficients of $\frac{(x-c)^{k}}{k!}$ in $p(x)$ are denoted $f_{k}(c)$. We always assume $f_{0}(c)=f(c)$, so that Peano differentiability of order zero is continuity and of order one is ordinary differentiability. The relation
$f_{n}(c)\Rightarrow Rf^{(n)}(c)$
follows from Taylor’s theorem and the structure of (1.1).
For $n\geq 2$, the converse implication
$Rf^{(n)}(c)\Rightarrow f^{(n)}(c)$
can fail everywhere for non-measurable functions, and even on large sets for measurable functions. We present an example illustrating the first case and show that the second arises from the failure of the converse to a generalized form of Taylor’s theorem.
For each nowhere dense closed set $H\subset\mathbb{R}$ and $n\geq 2$, there exists a function $f$ such that $f_{n}$ exists everywhere but $f^{(n)}(c)$ fails precisely on $H$. Thus, even for measurable functions, the implication $Rf^{(n)}(c)\Rightarrow f^{(n)}(c)$ fails dramatically. This naturally raises the question: does $Rf^{(n)}(c)$ imply $f_{n}(c)$?
A theorem of Marcinkiewicz and Zygmund affirms this implication under the assumption of measurability: if $f$ is measurable on $\mathbb{R}$ and $Rf^{(n)}(c)$ exists on a measurable set $E$, then $f_{n}(c)$ exists almost everywhere on $E$. Their argument is based on a more local result: if $f_{0}(c),\dots,f_{n-1}(c)$ exist, and the Riemann derivative corresponding to the nodes $\{0,1,2,4,\dots,2^{n-1}\}$ exists at $c$, then $f_{n}(c)$ exists.
This leads to the main focus of the present paper: we investigate which Riemann differences possess the property that
$f_{k}(c)\text{ exists for }0\leq k\leq n-1\text{ and }Rf^{(n)}(c)\text{ exists}\Rightarrow f_{n}(c)\text{ exists}.$
We call such differences those with the Marcinkiewicz-Zygmund (MZ) property. If this implication fails for a given $R$, then we can construct a function $g$ such that $g_{k}(0)=0$ for $k<n$, $R^{(n)}g(0)=0$, but $g_{n}(0)$ does not exist. Thus, the MZ property is equivalent to the implication
$g(h)=o(h^{n-1})\text{ and }Rg(h)=o(h^{n})\Rightarrow g(h)=o(h^{n}),$
where $Rg(h):=\sum_{k}a_{k}g(b_{k}h)$ is a generalized difference operator.
Previous results have shown that Riemann differences with geometric nodes of the form $\{0,1,q,\dots,q^{n-1}\}$ or $\{1,q,\dots,q^{n}\}$ for $q\notin\{-1,0,1\}$ satisfy the MZ property. This led to the conjecture that only such geometric-node differences could have the MZ property. However, this conjecture was disproven by the third-order example with nodes $\{-1,0,1,2\}$, which also satisfies the property (see [5]).
In this paper, we establish a complete classification of all Riemann differences with geometric nodes that possess the MZ property. We introduce a general analytic framework based on the recurrence
$D(h)=R(qh)-AR(h),$
and determine precisely when such a relation, combined with decay conditions $D(h)=o(h^{n})$ and $R(h)=o(h^{n-1})$, forces $R(h)=o(h^{n})$. Our main result shows that this happens if and only if $|A|$ lies outside a critical annulus determined by $q$ and $n$. The classification applies uniformly across the cases $|q|>1$ and $10<|q|<$ 1, and leads to a necessary and sufficient condition for the MZ property in terms of the roots of the characteristic polynomial of the difference.
We also demonstrate how this framework resolves various previously open examples, including several fourth-order Riemann differences that do satisfy the MZ property despite not conforming to earlier geometric templates. The techniques extend to asymmetric and shifted configurations and show that the MZ property is governed by the root geometry rather than the symmetry or placement of nodes.
Finally, we present an application illustrating how continuity at the origin can enlarge the class of first-order Riemann differences equivalent to the ordinary derivative. In particular, we show that continuity can restore differentiability in cases where, without it, the classical implication fails (cf. [1]).
2. Necessary conditions
In the introduction, we observed that the existence of a Riemann derivative does not imply the existence of the ordinary derivative, and that this failure can persist even when the function is measurable. We then asked a more refined question: under what conditions does the existence of a Riemann derivative imply the existence of the corresponding Peano derivative? This leads to the Marcinkiewicz-Zygmund property, which asserts that such an implication holds under appropriate regularity assumptions.
In this section, we present examples showing that the Peano implication can fail when those assumptions are violated. These examples demonstrate that continuity, measurability, and the existence of lower-order Peano derivatives are necessary for the MZ property to hold. They also motivate the formal framework we will use to classify which differences possess this property.
Marcinkiewicz and Zygmund proved that if $f$ is measurable on $\mathbb{R}$ and the Riemann derivative $Rf^{(n)}(x)$ exists on a measurable set $E$, then the $n$-th Peano derivative $f_{n}(x)$ exists almost everywhere on $E$. In particular, if $f$ is Schwartz differentiable on $\mathbb{R}$—i.e., if the symmetric difference quotient converges—then $f$ is also ordinarily differentiable almost everywhere. Remarkably, this conclusion holds even without assuming measurability of $f$. However, as the following example demonstrates, the measurability hypothesis becomes essential when $n\geq 2$.
Let $n\geq 2$ and fix a set of nodes $\{b_{1},b_{2},\ldots,b_{n+1}\}$. We will construct a non-linear solution of the Cauchy functional equation
(2.1) $f(x+y)=f(x)+f(y)$
such that $Rf^{(n)}(x)=0$ for all $x\in\mathbb{R}$. By a result of Darboux, such a function is nowhere continuous and, in particular, is not Peano differentiable at any point. That such a solution is also non-measurable was first established independently by Banach and Sierpiński in 1920.
Example 2.1.
Let $F=\mathbb{Q}(b_{1},b_{2},\ldots,b_{n+1})$ be the smallest subfield of $\mathbb{R}$ containing $\mathbb{Q}$ and the nodes. Since $F$ is countable, we can choose real numbers $\alpha$ and $\beta$ such that $\alpha/\beta\notin F$. Then $\{\alpha,\beta\}$ is linearly independent over $F$, and we may extend it to a Hamel basis $H$ of $\mathbb{R}$ over $F$.
Every $x\in\mathbb{R}$ has a unique finite representation
$x=d(x)\alpha+\sum_{i}d_{i}h_{i},$
where $d(x),d_{i}\in F$ and $h_{i}\in H$. Define $f(x)=d(x)$, the coefficient of $\alpha$ in the Hamel expansion of $x$. Then $f$ satisfies (2.1), since for any $x,y\in\mathbb{R}$,
$f(x+y)=d(x+y)=d(x)+d(y)=f(x)+f(y).$
Because $f(\alpha)=1$ and $f(\beta)=0$, $f$ is not linear (indeed $f(0)=0$), and hence $f$ is nowhere continuous and non-measurable.
For any $c\in\mathbb{R}$ and $h\in\mathbb{R}$, we have $d(c+b_{k}h)=d(c)+b_{k}d(h)$, so that
$\sum a_{k}f(c+b_{k}h)=d(c)\sum a_{k}+d(h)\sum a_{k}b_{k}=0$
by the system (1.1). Hence $Rf^{(n)}(c)=0$ for all $c\in\mathbb{R}$.
This function shows that the implication $Rf^{(n)}(c)\Rightarrow f_{n}(c)$ can fail everywhere when $f$ is not continuous at $c$. Thus, for the implication to hold, it is necessary that $f$ be continuous at $c$. More generally, as the next example illustrates, the existence of the first $n-1$ Peano derivatives at $c$ is also necessary.
Example 2.2.
Let $n\geq 2$, let $\{b_{1},\ldots,b_{n+1}\}$ be a set of nodes, and let $F$ be the field from Example 2.1. Define
$f(c+h)=\begin{cases}h^{n-1}&\text{if }h\in F,\\ 0&\text{otherwise}.\end{cases}$
Then $f$ is measurable and satisfies $f(c+h)=o(h^{n-2})$ as $h\to 0$. By the conditions (1.1), we again have $Rf^{(n)}(c)=0$. However, $f_{n-1}(c)=(n-1)!$ on $F$ and 00 on its complement, so $f_{n}(c)$ fails to exist.
We now formalize the obstruction to the MZ property. Suppose there exists a function $f$ and a point $c$ such that $f_{k}(c)$ exists for $k=0,\dots,n-1$, and $Rf^{(n)}(c)$ exists, but $f_{n}(c)$ does not. Define a function
$g(h)=f(c+h)-\sum_{k=0}^{n-1}\frac{f_{k}(c)}{k!}h^{k}-\frac{Rf^{(n)}f(c)}{n!}h^{n}.$
Then $g_{k}(0)=0$ for all $k<n$, and $R^{(n)}g(0)=0$ by the system (1.1). Moreover, $g_{n}(0)$ exists if and only if $f_{n}(c)$ exists. Thus, if $R$ is defined by
$Rg(h)=\sum_{k}a_{k}g(b_{k}h),$
then we may rephrase the MZ property as follows:
The difference $R$ has the Marcinkiewicz-Zygmund (MZ) property if, for all functions $g$, the conditions $g(h)=o(h^{n-1})$ and $Rg(h)=o(h^{n})$ imply $g(h)=o(h^{n})$.
Notation. In what follows, we distinguish between two related uses of the symbol $R$. When we write $Rf(h)$, we mean the application of a linear difference operator $R$, defined by
$Rf(h):=\sum_{k}a_{k}f(b_{k}h),$
where $\{a_{k},b_{k}\}$ are fixed coefficients and nodes. If these coefficients satisfy the moment system (1.1), then $R$ is called a Riemann difference of order $n$.
More generally, we say that a difference operator $R$ is of order $n$ if it annihilates all polynomials of degree less than $n$, but acts nontrivially on $x^{n}$:
$Rf(h)=0\quad\text{for all polynomials }f(x)=x^{m}\text{ with }0\leq m\leq n-1,$
$\quad\text{and}\quad R(x^{n})(h)=ch^{n}\text{ with }c\neq 0.$
In contrast, when we write $R(h)$, we refer to an abstract complex-valued function of $h$, which may or may not arise from applying a difference operator to a function. Theorems involving recursions like $D(h)=R(qh)-AR(h)$ operate at this more general level, where both $R(h)$ and $D(h)$ are understood as functions satisfying asymptotic and recurrence properties. This distinction allows us to formulate results in maximum generality while still applying them to concrete difference operators.
Normalization. When classifying differences up to the MZ property, it is convenient to reduce to a standard form. The following lemma justifies this reduction.
Lemma 2.3.
Let $R$ be a difference of order $n$ with coefficients $a_{k}$ and nodes $b_{k}$, and let $c\neq 0$, $r\neq 0$. Define the rescaled difference $R^{\prime}$ by
$R^{\prime}f(h):=\sum_{k}ra_{k}f\left(\frac{b_{k}}{c}h\right).$
Then $R$ has the MZ property if and only if $R^{\prime}$ does.
Proof.
The change of variables $h\mapsto ch$ implies $f(h)=o(h^{m})$ if and only if $f(ch)=o(h^{m})$. Rescaling the coefficients by $r$ does not affect asymptotic decay, so the MZ property is preserved. ∎
As a consequence, we may assume without loss of generality that one of the nodes is equal to $1$, and that one of the coefficients is equal to $1$.
Convention. Throughout this paper, we write $R(h)=o(h^{n})$ to mean that
$\lim_{h\to 0}\frac{|R(h)|}{|h|^{n}}=0,$
where $R(h)$ is complex-valued and defined on $\mathbb{R}\setminus\{0\}$. Similarly, if $Rf(h)=\sum a_{k}f(b_{k}h)$ is a difference operator applied to a function $f$, we say $Rf(h)=o(h^{n})$ if the same asymptotic relation holds.
3. Recurrence criteria for asymptotic decay
In this section, we develop a general analytic framework to determine when a function $R(h)\colon\mathbb{R}\to\mathbb{C}$, satisfying a recurrence of the form
$D(h)=R(qh)-AR(h),$
exhibits higher-order vanishing at zero. This setting arises naturally in the study of the Marcinkiewicz-Zygmund property for Riemann differences, where $R(h)$ corresponds to a difference operator applied to a function and $D(h)$ represents its increment.
Our goal is to determine when the conditions
$|D(h)|=o(h^{n})\quad\text{and}\quad|R(h)|=o(h^{n-1})$
force the improved decay $|R(h)|=o(h^{n})$. We show that this occurs if and only if the recurrence coefficient $A$ lies outside a critical annulus bounded by $|q|^{n-1}$ and $|q|^{n}$. The results apply to all real numbers $q\neq 0,\pm 1$.
Overview.
  • In the low-growth and high-growth regimes, we prove that $R(h)=o(h^{n})$ necessarily holds.
  • In the intermediate range $|q|^{n-1}<|A|\leq|q|^{n}$, we construct explicit counterexamples where $R(h)=o(h^{n-1})$ and $D(h)=o(h^{n})$ but $R(h)\not=o(h^{n})$.
  • The low-growth and failure cases are relatively straightforward, but the high-growth case requires delicate control of recurrence growth and asymptotic tracking.
We will provide the proofs only for the cases $|q|>1$, since the case $|q|<1$ follows from this one by the substitution
$R^{\prime}(h):=R(qh),\quad D^{\prime}(h):=\frac{-1}{A}D(h),$
under which the recurrence $D(h)=R(qh)-AR(h)$ is transformed into the equivalent relation
$D^{\prime}(h)=R^{\prime}\left(\frac{1}{q}h\right)-\frac{1}{A}R^{\prime}(h).$
3.1. Decay when $A$ lies outside the critical annulus
Theorem 3.1 (Low-growth case).
Let $|q|>1$, and let $D(h)$, $R(h)$ be complex-valued functions such that
$|D(h)|=o(h^{n}),\quad|R(h)|=o(h^{n-1}),$
and
$D(h)=R(qh)-AR(h),$
for some $A\in\mathbb{C}$ with $|A|\leq|q|^{n-1}$. Then $R(h)=o(h^{n})$.
Proof.
Fix $\epsilon>0$. Since $D(h)=o(h^{n})$, there exists $\delta>0$ such that $|D(h)|<\epsilon|h|^{n}$ whenever $0<|h|<\delta$. Fix such an $h\in(0,\delta)$, and define the sequence $h_{k}:=h/q^{k}$ for $k\in\mathbb{N}$.
By the recurrence relation,
$R(h_{k})=\frac{R(h_{k-1})-D(h_{k})}{A},$
we can expand $R(h)$ as a telescoping sum: